OpenStudy (anonymous):

i have 2 questions that involve fractions if someone could help.

OpenStudy (sam141101):

what are the questions?

OpenStudy (anonymous):

\[x/4 = 15\]

OpenStudy (sam141101):

sorry i suck at fractions

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you can solve this two ways one is to think, the other is the math teacher way the think way is to thing of a number that, when divided by 4, gives 15 the math teacher way is to "multiply both sides by 4" so \[\frac{x}{4}=15\\x=15\times 4\]

OpenStudy (anonymous):

so x = 60?

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

the check is easy \[\frac{60}{4}=15\]

OpenStudy (anonymous):

ok thanks. can you help me with one more? @misty1212

OpenStudy (misty1212):

sure

OpenStudy (anonymous):

my other one is similar. it's x/7 = 5, so would that be 35?

OpenStudy (misty1212):

eys

OpenStudy (misty1212):

yes

OpenStudy (anonymous):

thank you!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]